### v3

# LDPKiT: Superimposing Remote Queries for Privacy-Preserving Local Model Training

Abstract
Users of modern Machine Learning (ML) cloud services face a privacy conundrum—on one hand, they may have concerns about sending private data to the service for inference, but on the other hand, for specialized models, there may be no alternative but to use the proprietary model of the ML service. In this work, we present LDPKiT, a framework for non-adversarial, privacy-preserving model extraction that leverages a user’s private in-distribution data while bounding privacy leakage. LDPKiT introduces a novel superimposition technique that generates approximately in-distribution samples, enabling effective knowledge transfer under local differential privacy (LDP). Experiments on Fashion-MNIST, SVHN, and PathMNIST demonstrate that LDPKiT consistently improves utility while maintaining privacy, with benefits that become more pronounced at stronger noise levels. For example, on SVHN, LDPKiT achieves nearly the same inference accuracy at $\epsilon=1.25$ as at $\epsilon=2.0$, yielding stronger privacy guarantees with less than a 2% accuracy reduction. We further conduct sensitivity analyses to examine the effect of dataset size on performance and provide a systematic analysis of latent space representations, offering theoretical insights into the accuracy gains of LDPKiT.
I Introduction
In specialized domains such as medical imaging and cybersecurity, users who wish to employ machine learning (ML) solutions may have no choice but to use an ML model trained by a service provider, such as IBM Watson Health and Darktrace [1, 2, 3]. Because of the lack of publicly available datasets in such privacy-sensitive areas, organizations that have or can acquire access to large proprietary datasets are able to train ML models that average users or the open-source community simply cannot replicate. As a result, when using such a model, a user may have no choice but to send their inputs to the ML service provider for inference, opening themselves to a loss of privacy.
Unfortunately, many ML service providers do not prioritize the privacy of their users. In fact, some service providers explicitly state in their privacy policies that user-provided inference inputs may be stored indefinitely for purposes such as model training and advertising. Additionally, they may share these inputs with third parties to facilitate external services. [4, 5, 6, 7]. Such data practices can directly violate privacy regulations like the European Union’s General Data Protection Regulation (GDPR) [8]. Furthermore, several cases have emerged where providers have misused user data [9, 10, 11, 12]. In 2016, for example, Yahoo secretly colluded with the US government’s digital communication surveillance and used its custom spam and child pornography detection system to monitor user emails [9, 10, 11]. In 2022, the Amazon Ring Doorbell was found to have disclosed users’ video and audio footage to the police without authorization [12]. Even the recent large language models (LLMs) have incurred data leakage [13, 14], leading to many organizations banning their use.
Various methods for protecting privacy during inference have been proposed. For instance, homomorphic encryption [15, 16] and hardware-enforced trusted execution environments [17] ensure that the user’s inputs are always encrypted during inference, so that the service provider cannot see their contents. However, they cannot protect against a malicious or faulty model that leaks information about the inputs. Moreover, they require the cooperation of the ML service provider to deploy such privacy-preserving measures.
The strongest privacy guarantee a user can obtain is to perform model inference locally, so that the user’s private data never leaves their control. However, in domains where training data is scarce and open-source models do not exist, this is not possible. Moreover, the current business model of many ML companies is to deliberately keep their models, code and training data proprietary, so that they may maintain a competitive advantage [18, 19, 20]. They argue that proprietary models support the significant investment they make in improving their models, including acquiring high-quality proprietary datasets for training [21, 22]. As a result, it is also not possible for users to obtain proprietary models from ML service providers to run locally. In such circumstances, users may resort to model extraction: performing queries on the ML service provider’s model and using the results to train their own local model. While ML service providers generally do not condone model extraction, it is difficult for them to stop, and similar techniques, such as result caching to improve edge-device performance, are widely adopted as standard optimizations [23]. Moreover, model extraction still requires users to perform and pay for queries made to the ML service provider, so the provider may tolerate model extraction so long as their business model remains viable.
Motivated by these observations, we propose a model extraction-inspired privacy-preserving technique, LDPKiT (Local Differentially-Private and Utility-Preserving Inference via Knowledge Transfer), which enables users to perform a privacy-protective model extraction on specialized models. However, effective extraction requires users’ queries to be representative of the model’s training data distribution, so ideally, they should use real data. Due to the scarcity of public data in privacy-sensitive domains, users must use their own real data for extraction. Thus, LDPKiT uses local differential privacy (LDP) to bound privacy leakage by injecting random noise into inference queries. Further, users will likely only have a small amount of unlabelled in-distribution data; otherwise, they would be able to train a model themselves. LDPKiT addresses this issue with a novel data augmentation technique that superimposes noised input pairs, generating more in-distribution queries for model extraction. Such a setting naturally arises in practice, for example, in small hospitals with limited but sensitive in-distribution medical scans, or in financial institutions with private in-distribution transaction records. In these scenarios, LDPKiT provides a mechanism for leveraging powerful remote models while ensuring that sensitive user data remains protected.
We acknowledge that model extraction is controversial. In an extreme case, an adversary that extracts an ML service provider’s commercial model with the intention to resell that model and compete with the ML service provider (i.e., model theft) is considered adversarial and may be in direct violation of the provider’s terms of service. We naturally do not advocate such uses of model extraction. In contrast, we focus on privacy protection and employ extraction techniques in a non-adversarial way. Our approach differs from competitive model theft in several important ways: 1) LDPKiT’s extracted model is not competitive with the provider’s model due to the addition of LDP noise and would not violate the terms of service of most commercial model providers [24]. 2) Users perform a large number of queries to the model provider and presumably compensate the model provider for these queries. Since preventing model extraction, especially for small models, is inherently challenging, it is often accounted for in query pricing structures [25, 26, 27].
Our analysis is guided by the following research questions:
RQ1. Does LDPKiT recover the utility impacted by LDP noise? (Section IV-B)
RQ2. Why superimposition works in LDPKiT? (Section IV-C)
RQ3. How do the size of the private dataset and the number of queries impact LDPKiT? (Section IV-D)
Contributions. LDPKiT generates inference datasets that are both privacy-protective and close to the distribution of sensitive datasets, allowing users to accurately annotate their private data with bounded privacy leakage. We evaluate LDPKiT across diverse applications and multiple model architectures, systematically examining its performance and sensitivity to dataset sizes. Additionally, we conduct in-depth analyses of latent space representations to uncover factors contributing to LDPKiT’s effectiveness.
II Overview
Figure 1: Comparison of accuracies on SVHN and Fashion-MNIST: Extracting a local model (ResNet-18) with OOD public data samples versus querying the remote model (ResNet-152) with $\epsilon$-LDP protected private data.
II-A Motivation
Our scenario addresses the use of ML services for labelling privacy-sensitive data. We aim to protect the user’s data when querying a service provider’s model during inference, where the untrusted service provider may maliciously process user inputs without consent, violating regulations such as GDPR. To mitigate privacy risks, users without access to sufficient labelled data with which to train their models may opt to extract the service provider’s model. However, successful extraction requires in-distribution data, and relevant public data is usually scarce or unavailable in privacy-sensitive applications. Extracting a model with out-of-distribution (OOD) data fails to produce a functional model. We illustrate this with a simple experiment: we attempt to extract a ResNet-18 model trained on the Fashion-MNIST dataset using samples from SVHN, and, conversely, extract a model trained on the SVHN dataset using samples from Fashion-MNIST. As shown in the two leftmost bars of each group in Figure 1, the resulting accuracies on both datasets are close to random guessing (approximately 10% for these 10-class datasets). In contrast, extracting a model with in-distribution samples, even with noise added to those samples, results in higher accuracy. Specifically, ResNet-18 models trained on noisy in-distribution data, modified using $\epsilon$-LDP with strong privacy guarantees (i.e., low $\epsilon$ values), achieve an average accuracy of around 20%. Although 20% accuracy does not represent a successful model extraction, this result motivates the adoption of $\epsilon$-LDP as the foundation for our privacy-preserving framework. We subsequently enhance this baseline with techniques designed to recover the accuracy degradation incurred from LDP noise. These enhancements yield a fourfold improvement in accuracy under the same privacy level.
II-B Preliminaries and General Setup
There are two parties involved in our scenario: a model provider with the remote model and a user with a small set of unlabelled sensitive data. We use $\mathcal{M}_{\rm R}$}} to denote the remote model and $\mathcal{M}_{\rm L}$}} to denote the model extracted by the user. $\mathcal{D}_{\rm priv}$} denotes the privacy-sensitive dataset the user owns and wishes to label and protect, and $|{\mathcal{D}_{\rm priv}}|$ represents the number of samples in $\mathcal{D}_{\rm priv}$}. To protect $\mathcal{D}_{\rm priv}$}, we add $\epsilon$-LDP noise to each sample, resulting in $\mathcal{D}_{\rm protected}$} (and naturally $|{\mathcal{D}_{\rm priv}}|=|{\mathcal{D}_{\rm protected}}|$). $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm protected}}|$ are generally expected to be small. We thus propose to augment the number of privacy-protected samples available for model extraction, creating an expanded inference dataset, $\mathcal{D}_{\rm infer}$}. The size of $\mathcal{D}_{\rm infer}$} ($|{\mathcal{D}_{\rm infer}}|$) is determined by the user’s query budget, i.e., the number of queries the user is willing to make to $\mathcal{M}_{\rm R}$}, given the pricing model of the service. This allows users to flexibly select a subset or the entire set of augmented samples generated from $\mathcal{D}_{\rm protected}$}. These augmented samples are labelled by querying $\mathcal{M}_{\rm R}$} and subsequently used to train $\mathcal{M}_{\rm L}$}. We use SIDP to denote a baseline use of LDP that does not involve model extraction. In SIDP, the user simply queries $\mathcal{M}_{\rm R}$} with $\mathcal{D}_{\rm protected}$} and accepts the returned results as correct, despite potential errors introduced by the LDP noise in $\mathcal{D}_{\rm protected}$}. The two rightmost bars in Figure 1 show the performance of $\mathcal{M}_{\rm R}$} a user can expect when using SIDP. Currently, LDPKiT supports only the image modality, where superimposition maps naturally to pixel-level operations. Extending this concept to text or other data types would require new mechanisms (e.g., embedding-level or token-level mixing), which we leave as future work. To improve the prediction accuracy on $\mathcal{D}_{\rm priv}$}, LDPKiT proposes to extract an $\mathcal{M}_{\rm L}$} with samples from $\mathcal{D}_{\rm infer}$} generated using two data augmentation methods: LDPKiT-Rand and LDPKiT-Sup, which we discuss in Section III-B.
II-C Threat Model
The user’s goal is to label samples in $\mathcal{D}_{\rm priv}$} with reasonable accuracy when interacting with an ML service in a privacy-preserving way. We assume that $\mathcal{M}_{\rm R}$} is honest but curious. It honestly answers the user’s queries but may record the queries and results to infer information about the user. To mitigate this threat, the user applies $\epsilon$-LDP to their data before querying $\mathcal{M}_{\rm R}$}, limiting the amount of information that can be inferred from individual samples. The user knows $\mathcal{M}_{\rm R}$}’s task and functionality, but has no access to its underlying model architecture and dataset on which it was trained. $\mathcal{M}_{\rm L}$} aims to learn $\mathcal{M}_{\rm R}$}’s functionality, relying only on the black-box access limited to hard label responses from query interactions.
II-D Privacy Guarantee
LDPKiT ensures the privacy of $\mathcal{D}_{\rm priv}$} during inference by applying LDP noise to each data sample. This section discusses the key concepts and methodologies essential for designing effective privacy-preserving data augmentation mechanisms.
Definition II.1.
$\epsilon$}-Local Differential Privacy (LDP). We define $\epsilon$}-LDP as follows [28]: A randomized algorithm $\mathcal{A}$ satisfies $\epsilon$}-LDP if for all pairs of values and all sets $\mathcal{S}$ of possible outputs, where $\mathcal{S}\subseteq\text{Range}(\mathcal{A})$,
$\Pr[\mathcal{A}(v_{1})\in\mathcal{S}]\leq e^{\epsilon}\Pr[\mathcal{A}(v_{2})\in\mathcal{S}]$ (1)
A lower $\epsilon$ value indicates a tighter bound of the equation and a stronger privacy guarantee. We choose $\epsilon$ values primarily based on the industry standards [29, 30] and empirically verify that privacy-protected data is resilient to data reconstruction attacks [31]. We apply Laplacian noise to satisfy the $\epsilon$-LDP privacy guarantee, and the definition is as follows.
Definition II.2.
Laplacian Mechanism. The Laplacian mechanism of LDP adds noise drawn from the Laplacian distribution, with the probability density function (PDF) defined as follows for a variable $z$ and a scaling factor $\lambda\propto\frac{1}{\epsilon}$:
$L(z,\lambda)=\frac{1}{2\lambda}\exp\left(-\frac{|z|}{\lambda}\right)$ (2)
For the design of the candidate dataset used for inference, $\mathcal{D}_{\rm infer}$}, we also leverage LDP’s post-processing property to facilitate utility recovery. If an algorithm safeguards an individual’s privacy, a data analyst cannot compromise privacy—whether under the formal definition or in any intuitive sense—merely by reflecting on the algorithm’s output [32]. Post-processing is commonly used in DP schemes to improve the interpretability or accuracy of differentially-private data analysis [33]. The formal definition is as follows.
Definition II.3.
Post-Processing Property in $\epsilon$}-LDP. The post-processing property of LDP [32] states that if a randomized algorithm $\mathcal{A}$ satisfies $\epsilon$}-LDP, then for any deterministic or randomized function $g$ that does not depend on the original input values, the composed mechanism $g(\mathcal{A}(\cdot))$ also satisfies $\epsilon$}-LDP. Specifically, for all $v_{1},v_{2}$ (i.e., all values in the domain) and for all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$:
$\Pr[g(\mathcal{A}(v_{1}))\in\mathcal{T}]\leq e^{\epsilon}\Pr[g(\mathcal{A}(v_{2}))\in\mathcal{T}].$ (3)
The post-processing property ensures that applying a dataset-independent function $g$ to the output of $\mathcal{A}$ cannot weaken the privacy guarantees provided by $\epsilon$}-LDP, as long as $g$ itself does not access the raw data.
Figure 2: LDPKiT system overview.
III Design
Figure 2 describes both SIDP and LDPKiT. LDPKiT has four main stages: I. post-processing on noise-injected $\mathcal{D}_{\rm protected}$} to form a larger inference dataset, $\mathcal{D}_{\rm infer}$}, II. remote model inference with post-processed privacy-protected data from $\mathcal{D}_{\rm infer}$}, III. local training with inference results from $\mathcal{M}_{\rm R}$} and $\mathcal{D}_{\rm infer}$}, and IV. execution of users’ private inference queries on $\mathcal{M}_{\rm L}$} with $\mathcal{D}_{\rm priv}$}. $\mathcal{D}_{\rm infer}$} is generated from $\mathcal{D}_{\rm protected}$} using LDPKiT, with its size controlled by a user-defined query budget, as each query to $\mathcal{M}_{\rm R}$} incurs a financial cost. To optimize query efficiency, methods such as active learning [34] and core-set selection [35] could be used to choose informative subsets from $\mathcal{D}_{\rm infer}$}. We leave the investigation of such optimization techniques to future work and, for now, we assume that the user randomly samples a subset from the maximum available $\mathcal{D}_{\rm infer}$} for querying. LDPKiT may also be applied to an online learning setting where the user can iterate through the entire process and periodically train $\mathcal{M}_{\rm L}$} using $\mathcal{M}_{\rm R}$}’s predictions on new inference queries. SIDP, in contrast, does not involve $\mathcal{M}_{\rm L}$} training. It queries $\mathcal{M}_{\rm R}$} directly with data points from $\mathcal{D}_{\rm protected}$} and accepts $\mathcal{M}_{\rm R}$}’s outputs directly. Since $|{\mathcal{D}_{\rm protected}}|<|{\mathcal{D}_{\rm infer}}|$, SIDP has a lower querying cost than LDPKiT, but at the cost of utility.
III-A Preliminary Experiments
[…]
To validate our idea, we conduct a preliminary experiment on SVHN. Specifically, we construct a relatively small $\mathcal{D}_{\rm priv}$} with 500 samples randomly selected from SVHN. We set ResNet-152 as $\mathcal{M}_{\rm R}$} and ResNet-18 as $\mathcal{M}_{\rm L}$}. To ensure reliable results, we average the accuracies over three runs and across three different dataset splits. We set $\epsilon=1.5$ so that $\mathcal{M}_{\rm R}$}’s inference accuracy under SIDP (with a single layer of noise applied to each data point) is only 14.16%. This configuration ensures that $\mathcal{M}_{\rm R}$} provides minimal yet sufficient information while preserving privacy. To evaluate the effectiveness of our two-layer data augmentation mechanism, we then apply an additional layer of random noise 499 times to each of the 500 samples in $\mathcal{D}_{\rm protected}$}, creating an augmented inference dataset, $\mathcal{D}_{\rm infer}$}, with 249,500 data points (i.e., $n=499$). Performing model extraction with $\mathcal{D}_{\rm infer}$} results in an $\mathcal{M}_{\rm L}$} with 34.07% accuracy, while performing extraction only with the 500 original samples from $\mathcal{D}_{\rm protected}$} results in an $\mathcal{M}_{\rm L}$} that essentially guesses randomly (i.e., 10% accuracy). Compared to querying directly using $\mathcal{D}_{\rm protected}$} (i.e., SIDP), which gets 14.16% accuracy, this provides a 20% increase with no loss of privacy. A dependent two-sample t-test result shows that the improvement is statistically significant ($p<0.05$).
However, this naïve random-noise-based utility recovery mechanism does not offer an optimal utility-privacy trade-off, not only due to the limited size of $\mathcal{D}_{\rm priv}$} (500 samples) but also the undirected nature of the random noise augmentation. We hypothesize that data distributions form structured latent spaces, where queries constructed closer to decision boundaries could provide more informative training signals. Motivated by this insight, we explore a superimposition-based augmentation approach that refines the post-processed LDP noise. This approach aims to help $\mathcal{M}_{\rm L}$} better capture the decision boundaries, ultimately achieving higher accuracy on $\mathcal{D}_{\rm priv}$}. Under the same experimental setup, ResNet-18 achieves 45.69% accuracy—a 10% improvement over the naïve random noise augmentation.
III-B LDPKiT’s Data Augmentation Mechanism
LDPKiT applies Laplacian noise to the privacy-sensitive data in $\mathcal{D}_{\rm priv}$} to generate $\mathcal{D}_{\rm infer}$} for remote model inference, ensuring the $\epsilon$-LDP guarantee [28]. Since $\epsilon$ is defined in the context of LDP, an equal amount of noise is added to each data sample in $\mathcal{D}_{\rm priv}$} based on the $\epsilon$ value. We assume that queries in $\mathcal{D}_{\rm priv}$} are independent and identically distributed (i.i.d.), and the formulation of $\epsilon$-LDP is valid for each individual data point in $\mathcal{D}_{\rm priv}$}. Consequently, the privacy leakage is non-cumulative and bounded by $\epsilon$ per query. Given a private dataset $\mathcal{D}_{\rm priv}$} containing $|{\mathcal{D}_{\rm priv}}|$ i.i.d. data points, our goal is to create an augmented privacy-preserving $\mathcal{D}_{\rm infer}$} dataset safeguarded by $\epsilon$-LDP for querying $\mathcal{M}_{\rm R}$} and training $\mathcal{M}_{\rm L}$}. We propose two methods for data augmentation: the first involves adding different random noise samples to each data sample in $\mathcal{D}_{\rm protected}$} as described in Section III-A. The second involves superimposing pairs of samples from $\mathcal{D}_{\rm protected}$}. We emphasize in both cases that since the data augmentation is performed on $\mathcal{D}_{\rm protected}$}, the privacy level of both augmentation techniques remains the same as the original $\mathcal{D}_{\rm protected}$} under the post-processing property of LDP.
LDPKiT-Rand. We formally describe the augmentation technique mentioned previously in Section III-A here: for each $d_{\text{priv}_{i}}\in{\mathcal{D}_{\rm priv}}$, we construct an augmented dataset of the form: $d_{\text{infer}_{i}}\in{\mathcal{D}_{\rm infer}}=d_{\text{priv}_{i}}+\mathcal{L}_{i}+\mathcal{L}_{j}$, where $\mathcal{L}_{i}$ and $\mathcal{L}_{j}$ are Laplacian noise samples drawn independently with the same scale. The first noise layer, $\mathcal{L}_{i}$, is used to generate the initial privacy-protected dataset: $d_{\text{protected}_{i}}=d_{\text{priv}_{i}}+\mathcal{L}_{i}$, where $i=1,2,\dots,|{\mathcal{D}_{\rm protected}}|$. We then apply the second noise layer, $\mathcal{L}_{j}$, where $j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|-1$, as post-processing to create an expanded augmented dataset for querying. While it is theoretically possible to generate an infinite number of such noisy variants per data point, we cap the size of the maximum possible augmented dataset $\mathcal{D}_{\rm infer}$} to $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ to make it comparable to LDPKiT-Sup below.
LDPKiT-Sup. LDPKiT-Rand introduces unrelated noise that may cause the resulting data to deviate further from the distribution of $\mathcal{D}_{\rm priv}$}. LDPKiT-Sup is thus proposed as an alternative to better preserve the characteristics of $\mathcal{D}_{\rm priv}$}. After applying the first layer of $\epsilon$-LDP random noise to each data point in $\mathcal{D}_{\rm priv}$} to produce $\mathcal{D}_{\rm protected}$}, we combine pairwise samples of $\mathcal{D}_{\rm protected}$} for all pairs by averaging corresponding pixel values. Each augmented inference data point $d_{\text{infer}_{i}}\in{\mathcal{D}_{\rm infer}}$ is computed as $d_{\text{infer}_{i}}=avg(d_{\text{protected}_{i}},d_{\text{protected}_{j}})\text{ where }i,j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|,i\neq j$. This process results in at most $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ permutations in the protected inference dataset $\mathcal{D}_{\rm infer}$}.
Post-processed data samples in the augmented inference dataset $\mathcal{D}_{\rm infer}$}, generated with LDPKiT-Rand and LDPKiT-Sup, adhere to $\epsilon$-LDP and satisfy the post-processing property of LDP in Definition II.3. When querying $\mathcal{M}_{\rm R}$} with $\mathcal{D}_{\rm infer}$}, the user has the option to utilize either the entire available augmented inference dataset or a subset of it. We analyze the effect of $|{\mathcal{D}_{\rm infer}}|$ in Section IV-D. In Appendix A-A, we prove that the post-processed noisy data points in the augmented inference dataset $\mathcal{D}_{\rm infer}$}, generated using both LDPKiT-Rand and LDPKiT-Sup, adhere to $\epsilon$-LDP and satisfy the post-processing property of LDP.
Figure 3: Example of a noised Fashion-MNIST data point with label 5 (sandal) and $\epsilon$ set to 1.5.
IV Evaluation
In this section, we discuss the evaluation results and answer our research questions in Section I with empirical analysis.
(a) SVHN
(b) Fashion-MNIST
(c) PathMNIST
Figure 4: Inference accuracy comparisons on $\mathcal{D}_{\rm priv}$} with various $\epsilon$: ResNet-152 ($\mathcal{M}_{\rm R}$})’s SIDP versus $\mathcal{M}_{\rm L}$} trained using LDPKiT-Rand and LDPKiT-Sup. The results show the rank of performance as LDPKiT-Sup > LDPKiT-Rand > SIDP
IV-A Experimental setup
We run our experiments on two machines. One has two GPUs, NVIDIA GeForce RTX 3090 and 4090, with 24GB of dedicated memory, and an Intel 12th Gen i7-12700 CPU with 12 cores and 64GB of RAM. The other has two NVIDIA GeForce RTX 4090 GPUs and an AMD Ryzen Threadripper PRO 5955WX CPU with 16 cores and 64GB of RAM. The underlying OS are 64-bit Ubuntu 22.04.3 LTS and Ubuntu 24.04 LTS, respectively. We use Python 3.9.7 and PyTorch v2.1.2 with CUDA 12.1.
We evaluate LDPKiT on three diverse datasets: SVHN [36], Fashion-MNIST [37] and PathMNIST from MedMNIST2D for medical imaging in pathology [38]. For ML models, we use ResNet-152 [39] as $\mathcal{M}_{\rm R}$}}, and ResNet-18 [39] and MobileNetV2 [40] as $\mathcal{M}_{\rm L}$}}. The $\mathcal{M}_{\rm L}$} models are initialized with random weights.
To allocate data points for $\mathcal{D}_{\rm priv}$}, instead of using the default training split, we train ResNet-152 ($\mathcal{M}_{\rm R}$}) on 35k, 48,257 and 89,996 data points for Fashion-MNIST, SVHN and PathMNIST, respectively. A holdout dataset, $\mathcal{D}_{\rm val}$}, is used to evaluate $\mathcal{M}_{\rm L}$}’s generalizability. For Fashion-MNIST, the candidate pool of $\mathcal{D}_{\rm priv}$} has 25k data points and $\mathcal{D}_{\rm val}$} has 10k. For SVHN, the candidate pool for $\mathcal{D}_{\rm priv}$} has 25k data points and $\mathcal{D}_{\rm val}$} has 26,032. For PathMNIST, the candidate pool for $\mathcal{D}_{\rm priv}$} has 10,004 data points and $\mathcal{D}_{\rm val}$} has 7,180. To simulate a scenario that aligns with our setting, where users have limited private data, we set $|{\mathcal{D}_{\rm priv}}|=1,500$ for experiments in Sections IV-B. Specifically, we construct $\mathcal{D}_{\rm priv}$} by randomly selecting 1,500 data points in a balanced manner from the candidate pool so that each class is equally represented (balanced). This balanced sampling helps approximate identical distribution conditions across classes, reducing potential biases introduced by class imbalance. The effect of varying sizes of $\mathcal{D}_{\rm infer}$} and $\mathcal{D}_{\rm priv}$} is studied in Section IV-D. All experiments are repeated over three random seeds on three random subset splits to determine the statistical significance of our findings. The results of the dependent two-sample t-test confirm that all improvements are statistically significant (i.e., $p<0.05$).
We report the baseline SIDP accuracy using the labels returned by $\mathcal{M}_{\rm R}$}} on $\mathcal{D}_{\rm protected}$}, while the LDPKiT accuracy reflects the accuracy of $\mathcal{M}_{\rm L}$} on $\mathcal{D}_{\rm priv}$}, where $\mathcal{M}_{\rm L}$} is trained using $\mathcal{M}_{\rm R}$}}’s labels on $\mathcal{D}_{\rm infer}$}. We follow the industry standards [29, 30] to set $\epsilon$ for each dataset and ensure that the SIDP accuracy of $\mathcal{M}_{\rm R}$} corresponds to $1.5-3\times$ random guessing (e.g., 15% to 30% accuracy for a 10-class dataset), allowing $\mathcal{M}_{\rm R}$} to provide minimal but meaningful information for training $\mathcal{M}_{\rm L}$}. Moreover, we make sure that the noise added is high enough so that LDPKiT-protected data is resilient to data reconstruction attacks (See Section IV-E). Specifically, we choose $\epsilon$ values of 2.0, 1.5, and 1.25 for SVHN, 2.0 and 1.5 for Fashion-MNIST, and 10.0 and 7.0 for PathMNIST. The $\epsilon$ choices are consistent among experiments of SIDP, LDPKiT-Rand and LDPKiT-Sup. Figure 3 shows a privacy-protected data sample in Fashion-MNIST’s $\mathcal{D}_{\rm infer}$} generated with LDPKiT-Sup and LDPKiT-Rand. Additional noisy samples are shown in Appendix A-B. Hyperparameter choices are discussed and reported in Appendix A-C.
Figure 5: Inference accuracy comparisons of LDPKiT-Rand on $\mathcal{D}_{\rm priv}$} with base $\epsilon=1.5$ and various post-processing $\epsilon$.
IV-B RQ1: LDPKiT’s Utility Recovery on $\mathcal{D}_{\rm priv}$}
In this section, we evaluate whether LDPKiT improves prediction accuracy on $\mathcal{D}_{\rm priv}$} compared to directly using $\mathcal{M}_{\rm R}$}’s returned labels on $\mathcal{D}_{\rm protected}$} (i.e., SIDP). To simulate a realistic scenario, we set $|{\mathcal{D}_{\rm priv}}|$ to 1,500, representing a practical data size a user might own. For variable control, we use the entire $\mathcal{D}_{\rm infer}$} with the size of $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ for querying $\mathcal{M}_{\rm R}$} and training $\mathcal{M}_{\rm L}$} to minimize variations from random subset splits.
When no privacy protection exists and thus no noise is added, $\mathcal{M}_{\rm R}$}’s average accuracies on noise-free $\mathcal{D}_{\rm priv}$} of Fashion-MNIST, SVHN and PathMNIST are 93.33%, 94.73% and 86.2%, respectively. The accuracies on $\mathcal{D}_{\rm val}$} are 93.22%, 96.30% and 81.24%, respectively.
To answer RQ1 quantitatively, we record the last epoch’s accuracies of $\mathcal{M}_{\rm L}$} on the original $\mathcal{D}_{\rm priv}$} and $\mathcal{M}_{\rm R}$}’s SIDP accuracies on $\mathcal{D}_{\rm protected}$}. We compare SIDP and LDPKiT in Figure 4. We also tabulate the numerical accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} in Tables VVI and VII in Appendix A-D1. The results demonstrate that $\mathcal{M}_{\rm L}$} with LDPKiT can almost always achieve higher prediction accuracies on $\mathcal{D}_{\rm priv}$} compared to SIDP (the leftmost blue bar). For instance, when $\epsilon=2.0$, ResNet-152 ($\mathcal{M}_{\rm R}$})’s SIDP accuracy on SVHN is only 21.07%, whereas both LDPKiT-Rand and LDPKiT-Sup can recover the inference accuracy of the trained ResNet-18 ($\mathcal{M}_{\rm L}$}) to approximately 85%. However, the improvement is not always significant using LDPKiT-Rand. When employing LDPKiT-Rand, ResNet-18 achieves accuracies of only 68.52% and 58.07% on Fashion-MNIST with $\epsilon$ values of 2.0 and 1.5, respectively, and MobileNetV2’s accuracies fall below 60%. On PathMNIST, ResNet-18’s performance is also inadequate, and MobileNetV2’s accuracies with LDPKiT-Rand are even lower than SIDP. The results match our expectation that data generated by simply adding another layer of random noise is not representative enough for effective training. In contrast, the evaluation results of LDPKiT-Sup reported in parallel are constantly better than LDPKiT-Rand.
From the LDPKiT-Rand and LDPKiT-Sup bars in Figure 4, we observe that the local models learn more effectively using $\mathcal{D}_{\rm infer}$} created by superimposing noised $\mathcal{D}_{\rm priv}$} data points, i.e., LDPKiT-Sup, rather than applying arbitrary random noise as in LDPKiT-Rand. For instance, LDPKiT-Sup improves ResNet-18’s accuracy to 81.78% on Fashion-MNIST when $\epsilon$ is 2.0, compared to 28.69% with SIDP and 68.52% with LDPKiT-Rand. Similarly, ResNet-18’s accuracy is recovered to 82.11% on PathMNIST when $\epsilon$ is 10.0, far exceeding the 28.74% SIDP and 43.90% LDPKiT-Rand accuracies. Notably, despite the low SIDP baseline accuracy (11.87%–21.07%), LDPKiT-Sup achieves comparable accuracy even under more stringent privacy settings, consistently reaching over 80% accuracy on SVHN with both ResNet-18 and MobileNetV2. This demonstrates that LDPKiT-Sup provides stronger privacy guarantees without sacrificing utility. Overall, the experimental results confirm that LDPKiT effectively recovers most of the utility loss caused by LDP noise under strictly privacy-preserving noise levels.
Note that the base noise and the post-processing noise in LDPKiT-Rand are set to the same noise level (i.e., $\epsilon$ value) by default, as described in Section III-B. While it may seem intuitive to adjust LDPKiT-Rand by lowering the noise level of the post-processing layer (i.e., increasing $\epsilon$ for the second layer of noise) to achieve higher accuracy, empirical results suggest that maintaining the same $\epsilon$ value for the post-processing noise as the base noise is more efficient and accuracy-preserving. To explore the optimal configuration of LDPKiT-Rand, we conduct experiments on SVHN using a ResNet-18 model ($\mathcal{M}_{\rm L}$}), fixing the base noise $\epsilon_{1}$ at 1.5 and varying the post-processing noise $\epsilon_{2}$. In Figure 5, the left-most green bar represents LDPKiT-Sup’s inference accuracy, consistently outperforming LDPKiT-Rand across all the $\epsilon$ combinations. The orange bar (the seventh from left in Figure 5) indicates LDPKiT-Rand’s accuracy when $\epsilon_{1}=\epsilon_{2}$. The results show that deviating the post-processing noise significantly from the base noise reduces LDPKiT-Rand’s utility (e.g., inference accuracy drops to random guessing when $\epsilon>5$).
(a) (0,0)
(b) (9,2)
(c) (2,2)
Figure 6: Latent space plots of Fashion-MNIST class triplets (C0-T-shirt/top , C2-pullover, C9-ankle boot) and privacy-protected noisy data clusters generated with LDPKiT-Rand and LDPKiT-Sup ($\epsilon=2.0$). The observation is that the data clusters generated with LDPKiT-Sup are constantly closer to the target clusters than the clusters generated with LDPKiT-Rand.
TABLE I: Euclidean distances between centroids of clusters shown in Figure 6 on Fashion-MNIST, demonstrating that LDPKiT-Sup almost always generate data distribution that is closer to the target distribution than LDPKiT-Rand.
Figure Strategy Class(es) $d(C_{\textbf{\text{N}}},C_{0})$ $d(C_{\textbf{\text{N}}},C_{2})$ $d(C_{\textbf{\text{N}}},C_{9})$
6(a) Sup (0,0) 1.9835 2.5191 3.2017
Rand 0 2.1195 2.4666 2.9169
6(b) Sup (9,2) 3.1652 2.3538 2.0524
Rand 9 3.0937 2.5595 1.9333
6(c) Sup (2,2) 2.9662 1.7351 2.9171
Rand 2 2.7084 1.9433 2.7115
$d(C_{\text{N}},C_{X})$ is the Euclidean distance between the Noisy and Class X cluster centroids.
TABLE II: Frequency with which LDPKiT-Rand is more divergent from the target class triplets than LDPKiT-Sup when measuring with KL Divergence Ratio, demonstrating that LDPKiT-Sup produces more meaningful data samples than LDPKiT-Rand.
Dataset
Number of triplets
with DR > 1.0
Total number
of triplets
Frequency
SVHN 49 72 68.06%
Fashion-MNIST 25 36 69.44%
PathMNIST 15 22 68.18%
DR refers to the KL Divergence Ratio introduced in Section IV-C Definition IV.2.
Frequency refers to how often LDPKiT-Rand’s cluster is further from the target triplet cluster than LDPKiT-Sup (i.e., LDPKiT-Sup’s cluster is closer to target distribution).
IV-C RQ2: Latent Space Analysis
In Section IV-B, we empirically demonstrate that LDPKiT enhances inference accuracy on $\mathcal{D}_{\rm priv}$} compared to SIDP, with LDPKiT-Sup consistently outperforming LDPKiT-Rand. However, it is not apparent why LDPKiT-Sup is better than LDPKiT-Rand qualitatively. We hypothesize that the synthetic data points generated by LDPKiT-Sup are closer to the distribution of the original data points from target classes compared to LDPKiT-Rand, enabling more effective knowledge transfer by generating data points with higher quality, which leads to potentially better extraction of the decision boundaries. To evaluate our hypothesis, we plot and analyze the latent space of $\mathcal{D}_{\rm infer}$}. To replicate our framework in latent space, we first train a variational auto-encoder (VAE) on the data samples with $\epsilon$-LDP noise from three random classes of the datasets. We pick $\epsilon$ values of 2.0 for Fashion-MNIST and SVHN, and 7.0 for PathMNIST. Other hyperparameter choices are discussed in Appendix A-C. Our training objective is to visualize three distinct clusters representing the original noise-free data samples corresponding to the three target classes. To achieve this, we replace the conventional reconstruction loss with the triplet margin loss. Using the trained VAE model, we further generate and visualize two additional clusters: noisy samples produced by LDPKiT-Sup based on two of the three classes, and noisy samples produced by LDPKiT-Rand based on one class. We then compute the Euclidean distances from these noisy clusters (i.e., LDPKiT-Rand and LDPKiT-Sup) to the clean class clusters. This analysis allows us to evaluate whether LDPKiT-Sup offers a learning advantage over LDPKiT-Rand, particularly from the perspective of the latent space representation.
Figure 6 presents an example of Fashion-MNIST’s latent space visualization. Clusters of Class 0, Class 2 and Class 9 represent the original (noise-free) data points from those target classes. For instance, the cluster superimposed on Class 9 and Class 2 contains noisy superimposed data points generated with LDPKiT-Sup from Classes 9 and 2. The cluster randomized on Class 0 shows noisy data points generated with LDPKiT-Rand on the target Class 0. The key observation is that the data distribution of the noisy cluster created with LDPKiT-Sup is closer to the distribution of the target class(es) than that of LDPKiT-Rand. For instance, LDPKiT-Sup’s cluster is closer to the cluster of the target class (e.g., Class 0 or Class 2) than LDPKiT-Rand’s cluster in Figure 6(a) and Figure 6(c). In Figure 6(b), LDPKiT-Sup’s cluster is more spread out and has more overlapping area with the clusters of Class 2 and Class 9, whereas LDPKiT-Rand’s cluster remains more static and less representative, sitting in the middle of the clusters of Class 0, 2 and 9. Additional latent space analyses for triplet examples across other datasets are provided in Appendix A-D4.
Table I provides the numerical Euclidean distances between the centroids of each cluster pairs from Figure 6. Here, the distance values further validate the observations from the plots. We note, however, that the Euclidean distance between centroids is not always an ideal measure of similarity between clusters. For example, as shown in Figure 6(b), the cluster generated by LDPKiT-Sup visually contains more overlapping data points with the Class 9 cluster compared to LDPKiT-Rand. However, since LDPKiT-Sup incorporates points from both Class 2 and Class 9, its distribution is more dispersed than LDPKiT-Rand, which only contains data points from Class 9. Consequently, the centroid of LDPKiT-Rand is slightly closer to the centroid of the Class 9 cluster.
Therefore, to gain a more generalized and comprehensive understanding of the latent space representations, we also evaluate the Divergence Ratio (DR) of LDPKiT-Rand’s cluster compared to LDPKiT-Sup’s cluster with respect to the target triplet cluster. To quantify the divergence between clusters, we employ Kullback-Leibler (KL) Divergence [41] defined as follows.
Definition IV.1.
Kullback-Leibler (KL) Divergence. The KL Divergence of two probability distributions $P$ and $Q$ is defined as:
$D_{\text{KL}}(P\,||\,Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$
In our case, we denote the combination of three target clusters as the target triplet cluster, $C_{T}$, consisting of all the data points in the triplet classes. The cluster generated by LDPKiT-Sup, denoted as $C_{S}$, consists of all the superimposed combinations of two target classes, while LDPKiT-Rand’s generated cluster, $C_{R}$, consists of noisy data points generated on all three triplet classes. For a fair comparison, $C_{S}$ and $C_{R}$ contain the same number of points.
Definition IV.2.
Divergence Ratio (DR) between $C_{S}$ and $C_{R}$. The divergence ratio measures how often $C_{S}$ is less divergent from $C_{T}$ than $C_{R}$ to $C_{T}$ in the context of the KL Divergence.
$\text{DR}(\bar{C}_{R},\bar{C}_{S})=\frac{D_{\text{KL}}(\bar{C}_{T}\,\|\,\bar{C}_{R})}{D_{\text{KL}}(\bar{C}_{T}\,\|\,\bar{C}_{S})}$
where $\bar{C}_{S}$, $\bar{C}_{R}$ and $\bar{C}_{T}$ are the normalized distributions of $C_{S}$, $C_{R}$ and $C_{T}$. $\text{DR}(\bar{C}_{R},\bar{C}_{S})>1.0$ means that $\bar{C}_{S}$ is less divergent from $\bar{C}_{T}$ than $\bar{C}_{R}$ in this triplet class setting (i.e., $C_{S}$ is a better representation of $C_{T}$ than $C_{R}$).
TABLE III: Sensitivity analysis of LDPKiT on different dataset sizes, $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$, with ResNet-18 ($\mathcal{M}_{\rm L}$}).
Dataset $\epsilon$ LDPKiT Accuracy on $\mathcal{D}_{\rm priv}$} (%)
$|{\mathcal{D}_{\rm priv}}|$ 125 250 500 1k 1.5k
$|{\mathcal{D}_{\rm infer}}|$ 15,500 15,500 62,250 15,500 62,250 250k 15,500 62,250 250k 500k 15,500 62,250 250k 500k
Fashion- MNIST 2.0 Rand 14.76 18.36 36.40 17.71 32.47 52.93 18.06 34.45 57.90 63.37 14.81 34.68 57.20 64.66
Sup 25.87 28.93 31.42 24.20 37.64 52.95 25.11 47.18 66.48 73.66 25.14 45.08 68.98 74.35
SVHN 1.5 Rand 8.80 11.40 6.90 9.40 6.93 34.07 10.45 9.85 35.33 51.31 10.43 10.03 32.81 48.59
Sup 9.07 9.73 12.67 10.13 15.33 45.69 10.20 19.97 60.10 69.08 10.09 14.31 62.10 76.88
Path- MNIST 10.0 Rand 24.36 24.80 29.87 22.89 32.78 44.40 24.48 30.81 50.40 43.37 23.56 29.00 43.73 54.39
Sup 41.07 41.42 47.51 44.00 54.98 69.93 40.93 48.49 73.33 76.97 42.99 57.20 74.14 76.79
(a) Fashion-MNIST ($\epsilon=2.0$)
(b) SVHN ($\epsilon=1.5$)
(c) PathMNIST ($\epsilon=10.0$)
Figure 7: Comparison of ResNet18 ($\mathcal{M}_{\rm L}$})’s accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} with $|{\mathcal{D}_{\rm priv}}|=1.5k$ and various $|{\mathcal{D}_{\rm infer}}|$. The results show that LDPKiT-Sup demonstrates higher accuracy when $|{\mathcal{D}_{\rm infer}}|$ is larger.
Nuances may occur between different triplets. To evaluate the average-case scenario, we collect statistics across 130 triplets from SVHN, Fashion-MNIST, and PathMNIST and present the results in Table II. The analysis indicates that in approximately 68.56% of all the triplet cases, LDPKiT-Sup generates clusters that are less divergent from the target triplet cluster compared to LDPKiT-Rand, measured by KL Divergence. The finding that LDPKiT-Rand tends to diverge from the target distribution more frequently than LDPKiT-Sup matches our previous observations of higher accuracy achieved by LDPKiT-Sup and the insights gained from the latent space analysis. However, it is worth noting that the relationship between accuracy improvement and KL divergence can be complicated and requires further statistical analysis.
IV-D RQ3: Sensitivity Analysis of the Impact of $|{\mathcal{D}_{\rm infer}}|$ and $|{\mathcal{D}_{\rm priv}}|$ on LDPKiT
To study the trade-off among privacy ($|{\mathcal{D}_{\rm priv}}|$), cost ($|{\mathcal{D}_{\rm infer}}|$) and accuracy more thoroughly, we investigate how many private data samples ($|{\mathcal{D}_{\rm priv}}|$) are needed for augmentation and how many queries ($|{\mathcal{D}_{\rm infer}}|$) are needed to train an $\mathcal{M}_{\rm L}$} that achieves a reasonable accuracy on $\mathcal{D}_{\rm priv}$}. As shown in Table III, we generate several possible $\mathcal{D}_{\rm infer}$} datasets from $\mathcal{D}_{\rm priv}$} with varying sizes to examine the effect of $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ on $\mathcal{M}_{\rm L}$}’s accuracy. For example, when $\mathcal{D}_{\rm priv}$} contains 125 points, the largest possible $\mathcal{D}_{\rm infer}$} can have 15,500 points (i.e., $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$). If $|{\mathcal{D}_{\rm priv}}|$ is increased to 250, $|{\mathcal{D}_{\rm infer}}|$ can still be set to 15,500 by random selection or 62,250 at maximum. For variable control in analyzing LDPKiT’s sensitivity to $|{\mathcal{D}_{\rm priv}}|$, an example setting can be fixing $|{\mathcal{D}_{\rm infer}}|$ at 15,500 and comparing $\mathcal{M}_{\rm L}$}’s accuracies when $|{\mathcal{D}_{\rm priv}}|$ is 125 versus when it is 250. Similarly, we can assess the impact of $|{\mathcal{D}_{\rm infer}}|$ by fixing $|{\mathcal{D}_{\rm priv}}|$ at 250 and comparing $\mathcal{M}_{\rm L}$}’s accuracy for $|{\mathcal{D}_{\rm infer}}|=15,500$ and $|{\mathcal{D}_{\rm infer}}|=62,250$. To identify trends, we exhaustively evaluate various pairs of $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ for $|{\mathcal{D}_{\rm priv}}|$ ranging from 125 to 1.5k and record ResNet-18 ($\mathcal{M}_{\rm L}$})’s accuracies in each scenario.
As shown in Table III, LDPKiT exhibits lower sensitivity to variations in $|{\mathcal{D}_{\rm priv}}|$, which means that the user can get similarly accurate predictions without using more private data (i.e., LDPKiT provides similar utility with less privacy risk). For instance, the accuracies in the cases with fixed $|{\mathcal{D}_{\rm infer}}|$ and varying $|{\mathcal{D}_{\rm priv}}|$ are similar across the three benchmarks, e.g., on Fashion-MNIST with $\epsilon=2.0$, LDPKiT-Sup’s maximum accuracy difference among different $|{\mathcal{D}_{\rm priv}}|$ settings is below 5% when $|{\mathcal{D}_{\rm infer}}|=15,500$ and is around 0.69% when $|{\mathcal{D}_{\rm infer}}|=500k$. It also suggests that as long as $|{\mathcal{D}_{\rm infer}}|$ is reasonable, the user can generate queries with a smaller $|{\mathcal{D}_{\rm priv}}|$, which further reduces privacy risks. There is one case, however ($|{\mathcal{D}_{\rm infer}}|=250k$), where increasing $|{\mathcal{D}_{\rm priv}}|$ improves $\mathcal{M}_{\rm L}$}’s accuracies. This is likely because a larger training set increases the probability of including more representative and diverse data points that benefit training. However, this improvement is not guaranteed.
$|{\mathcal{D}_{\rm infer}}|$ determines the cost and has more impact on $\mathcal{M}_{\rm L}$}’s accuracy. Although $\mathcal{D}_{\rm priv}$} is less critical in size, it ultimately limits the maximum size of $\mathcal{D}_{\rm infer}$} we can generate. For example, on PathMNIST with $\epsilon=10.0$, when $|{\mathcal{D}_{\rm priv}}|=1.5k$, LDPKiT-Rand’s accuracy range from 23.56% with $|{\mathcal{D}_{\rm infer}}|=15,500$ to 54.39% with $|{\mathcal{D}_{\rm infer}}|=500k$. Similarly, LDPKiT-Sup’s accuracy also improved from 42.99% to 76.79% when $|{\mathcal{D}_{\rm infer}}|$ increases. In contrast, $\mathcal{M}_{\rm L}$}’s accuracy is always around 10% for SVHN when $|{\mathcal{D}_{\rm infer}}|$ is only 15,500, regardless of $|{\mathcal{D}_{\rm priv}}|$. Therefore, having a sufficient number of data points in $\mathcal{D}_{\rm infer}$} is crucial.
For a more straightforward visualization of $|{\mathcal{D}_{\rm infer}}|$’s impact, we also present $\mathcal{M}_{\rm L}$}’s accuracies with a fixed setting of $|{\mathcal{D}_{\rm priv}}|=1.5k$ and varying $|{\mathcal{D}_{\rm infer}}|$ as bar graphs in Figure 7. Notably, a reasonably large $\mathcal{D}_{\rm infer}$} is required for LDPKiT-Sup to have an adequate accuracy (e.g., 250k for PathMNIST and 500k for Fashion-MNIST and SVHN). While increasing $|{\mathcal{D}_{\rm infer}}|$ (i.e., higher querying cost) enhances accuracy, the improvement becomes marginal once $|{\mathcal{D}_{\rm infer}}|$ exceeds 500k. For instance, ResNet-18 ($\mathcal{M}_{\rm L}$}) achieves accuracies of 74.35%, 78.88% and 81.78% on $\mathcal{D}_{\rm priv}$} of Fashion-MNIST when trained on $\mathcal{D}_{\rm infer}$} with $\epsilon$ set to 2.0 and $|{\mathcal{D}_{\rm infer}}|$ set to 500k, 1M and 2.2M, respectively. In other words, it is unnecessary to query all possible data points in $\mathcal{D}_{\rm infer}$}, as discussed in Section IV-B. The trained $\mathcal{M}_{\rm L}$} with LDPKiT-Sup can achieve satisfactory accuracies using only about 20% of $\mathcal{D}_{\rm infer}$} for querying. Appendix A-D2 presents the evaluation results for other $\epsilon$ values on the same benchmarks, which match the conclusions drawn from Figure 7.
It is important to note that the subsets of $\mathcal{D}_{\rm infer}$} used to train $\mathcal{M}_{\rm L}$} are randomly selected in this section of the experiments. Advanced methods, such as active learning and core-set strategies [42, 35], could be employed to identify a smaller yet representative subset of $\mathcal{D}_{\rm infer}$}, effectively capturing the dataset’s overall distribution to train the model. Exploring such data selection methods for $\mathcal{D}_{\rm infer}$} size reduction is left for future work.
Currently, there are no universal guidelines for estimating a precise lower bound on $\lvert{\mathcal{D}_{\rm priv}}\rvert$ that guarantees reliable performance. The effective lower bound depends on several factors, including (i) the complexity of the remote model and task (e.g., number of classes, class imbalance), (ii) the degree of noise introduced by the LDP mechanism, and (iii) the diversity and representativeness of the available private samples. In practice, an end user with limited data can gauge feasibility through small-scale pilot runs. For example, if $\lvert{\mathcal{D}_{\rm priv}}\rvert$ is sufficiently large, the user may expand $\mathcal{D}_{\rm infer}$} by 5%, 10%, or 20% and monitor the resulting accuracy trend. Developing systematic methods to estimate realistic thresholds for $\lvert{\mathcal{D}_{\rm priv}}\rvert$ remains an important direction for future research.
TABLE IV: Classification accuracy and image similarity under various $\epsilon$ settings. Similarity is computed between original and noisy images (before reconstruction), and between original and denoised images (after reconstruction). For reconstruction quality: $\uparrow$ indicates higher is better; $\downarrow$ indicates lower is better.
Dataset $\bm{\epsilon}$ Recon.
ResNet-152
Acc. (%)
$\uparrow$
MSE $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SVHN 2.0 Before 16.80 0.0681 0.1210 0.7591
After 14.20 0.1524 0.0825 0.7183
1.5 Before 15.30 0.0929 0.0850 0.7747
After 14.40 0.1436 0.0829 0.7301
1.25 Before 12.50 0.1098 0.0681 0.7826
After 13.00 0.1382 0.0810 0.7347
Fashion -MNIST 2.0 Before 16.10 0.0942 0.3869 0.5568
After 16.70 0.1257 0.3180 0.4492
1.5 Before 14.80 0.1123 0.3406 0.5928
After 16.30 0.1395 0.2953 0.5044
Path MNIST 10.0 Before 22.78 0.0132 0.4823 0.2237
After 23.89 0.0853 0.3870 0.2593
7.0 Before 17.56 0.0155 0.3936 0.2507
After 21.89 0.1012 0.3110 0.2805
IV-E Evaluation of Data Reconstruction Risks
To rigorously evaluate whether the amount of LDP noise added to each private image is sufficient, we audit the privacy guarantees of LDPKiT following the methodology of [31]. Specifically, we conduct data reconstruction attacks using diffusion models (i.e., DDPM [43]) that learned the image priors (i.e., trained on the original images). Our experiment spans three datasets: SVHN, Fashion-MNIST, and PathMNIST. For each dataset, we randomly select a subset of 1,000 images (900 for PathMNIST) processed with LDPKiT-Sup, and quantitatively measure the data reconstruction success with established theoretical bounds: mean squared error (MSE), VGG-based learned perceptual image patch similarity (LPIPS) [44, 45], and structural similarity index measure (SSIM) [46]. We report results across different $\epsilon$ values in Table IV. Our measurements show that the reconstructed images are generally more dissimilar from the original images after reconstruction (i.e., lower SSIM, higher LPIPS, and higher MSE), indicating that the reconstruction fails. We also show that a fully trained classifier (ResNet-152) is unable to classify the reconstructed images correctly. We show some sample reconstructed images in Figure 12 in Appendix A-D3.
V Related work
V-A Knowledge transfer techniques
Knowledge distillation is a knowledge transfer technique that distills a large teacher model into a smaller student model while preserving model performance [47, 48, 49, 50, 51, 52]. The conventional use case is model compression, enabling the model deployment in a resource-restricted environment. Knowledge can be transferred in different forms, such as logits, model parameters, intermediate layers’ activations or features, and their interrelationships [48]. Knowledge transfer also has adversarial applications. A model extraction attack is an adversarial example in which the attacker reproduces a model stealthily by stealing its parameters, decision boundaries, or functionalities. It demonstrates that an iterative query-based knowledge transfer process from a high-performance model can be performed via a prediction query interface [53, 54, 55]. Model extraction can be challenging without knowledge of the victim model’s training data distribution [56]. Successful model extractions with partial or zero knowledge of the victim model and training data [57, 54, 56, 58, 59] may require more information than hard labels. Related defenses [60, 61, 62, 63] and analyses [64, 65] are in active research. Rather than model compression or adversarial extraction for the purpose of theft, we incorporate extraction techniques for privacy preservation in a non-adversarial manner to recover the utility loss brought by LDP. Our mechanism still demonstrates the privacy-utility trade-off rather than creating a competitive surrogate model that violates ML services’ terms of use.
V-B Noise injection and differential privacy (DP)
DP can be used locally [66] or globally [67, 68, 69], and both provide provable privacy guarantees. Global differential privacy (GDP) shares original input data with a trusted data curator, which then applies noise to the aggregated data. In this case, the curator has access to the original sensitive data. To remove this point of trust, LDPKiT applies local differential privacy (LDP). In LDP, the data source (i.e., the user) adds noise to each individual query before data transmission to the ML service, so she has full control over privacy protection. The side effect is that the noise is aggregated on a server, so LDP methods usually provide lower utility than GDP at the same level of privacy protection (i.e., noise level) [70]. While we add noise to original inputs before offloading inference to the cloud, similar to [71], noise can also be injected into inference frameworks deployed on a split computation setting [72, 73, 74, 75], where the DNN is partitioned between the cloud and edge devices. These schemes involve a white-box model, so noise can be added to intermediate representations, which differs from our setting, where we assume the model provider is reluctant to share the model with the user.
V-C Other privacy protection techniquess
One class of privacy protection methods is data encryption with homomorphic algorithms, which suffers from high computational overheads [15, 16]. For experiments on SVHN, LDPKiT begins to demonstrate lower computational overhead when the number of queries is around 800. Specifically, full knowledge transfer with LDPKiT on 800 data points only takes 20.73 seconds, compared to Lancelot [76], which takes 22.984 seconds for one inference on the same dataset, and OpenFHE [77], which requires significantly more time at 484.184 seconds. Furthermore, LDPKiT performs knowledge transfer once, enabling subsequent inference from the same sensitive data distribution on the trained local model with negligible overhead. In contrast, encryption schemes continue to impose per-query costs on the remote model, resulting in higher cumulative costs when the user wants to label a reasonably large dataset. Hardware-assisted inference in Trusted Execution Environments (TEEs) is an alternative approach [17]. A TEE is a secure area within a processor that provides a safe environment for sensitive code execution, preventing unauthorized access. Slalom puts the computation in a TEE to address inference privacy on remote services [17]. However, Slalom does not protect against the risks of side-channel attacks. Since TEEs have access to the original data, privacy breaches can still happen if the attackers compromise the TEEs [78, 79, 80]. Side-channels are not a threat to LDPKiT, since it does not transmit the original data, and any privacy leakage is bounded by the LDP noise.
V-D Federated Learning (FL)
FL has emerged as a popular approach to address privacy concerns by enabling collaborative model training without directly sharing users’ raw data with a central server [81, 82, 83, 84, 85]. However, FL primarily protects users’ training data, which is an orthogonal concern to our setting, and it does not mitigate privacy risks during inference. Moreover, FL is vulnerable to gradient inversion attacks that reconstruct local training data from gradient updates, which necessitates extra privacy protection [86, 87, 88, 89]. Finally, FL assumes that users possess labelled local datasets, which is not applicable in our case, where the objective is to annotate unlabelled data. Consequently, concerns regarding potential data reconstruction attacks and the integrity of the server imply that relying solely on using shared models from service providers like FL is insufficiently secure and privacy-preserving; therefore, FL is not an ideal solution in our scenario.
VI Discussion, Limitations and Future Work
VI-A Ethical Use of Model Extraction Principles
While our approach involves knowledge transfer from a remote model to a local model, it differs from adversarial model extraction attacks [57, 54, 56, 58, 59]. LDPKiT ethically integrates model extraction principles to protect user inference data information, prioritizing data privacy over adversarial exploitation.
Existing defenses against model extraction attacks often lack generalizability [25, 27], and advanced techniques [26] can bypass them, risking economic losses for commercial model owners. However, this is not a concern in our case. LDPKiT’s extraction objective focuses solely on protecting data privacy, not economic gain. At fairly privacy-protective noise levels, such as when SIDP drops to approximately 10% to 25%, the $\mathcal{M}_{\rm L}$} model generated by LDPKiT exhibits an accuracy trade-off for privacy ranging from 10% to 30%, depending on the noise level, model, and dataset. Notably, $\mathcal{M}_{\rm L}$} is not competitive with $\mathcal{M}_{\rm R}$}; for instance, ResNet-18 ($\mathcal{M}_{\rm L}$}) shows an accuracy drop of 12.96% on Fashion-MNIST with LDPKiT-Sup at $\epsilon=2.0$. Given this lack of competitiveness, $\mathcal{M}_{\rm L}$} does not serve the objectives of adversarial model extraction and adheres to the non-competition terms of use outlined by major commercial model providers [24].
In this work, our definitions of non-competitiveness, and non-stealthiness for ethical and non-adversarial extraction are scoped to the research setting: (i) Non-competitiveness means that the extracted $\mathcal{M}_{\rm L}$} models do not approach the utility or economic value of their $\mathcal{M}_{\rm R}$} counterparts and therefore cannot substitute or rival them in any commercial sense. (ii) Non-stealthiness indicates that LDPKiT does not attempt to evade detection or disguise its operation; its goal is explicitly protective, rather than adversarial. These working definitions emphasize that LDPKiT is designed and evaluated within a research context, to advance user data protection, rather than challenging or undermining existing commercial ecosystems. We further acknowledge that the notions of non-competitiveness and non-stealthiness require careful interpretation and, in practice, would necessitate formal reviews by legal and compliance experts.
VI-B Limitations and Future Work
As discussed in Section III, our per-query $\epsilon$-LDP privacy guarantee relies on the assumption that each data point in $\mathcal{D}_{\rm priv}$} is i.i.d.. If the data points are not i.i.d., the privacy guarantees can be weakened due to their mutual information. In practice, many real-world applications leverage data de-duplication techniques to maintain or approximate the i.i.d. assumption. For instance, cloud storage services such as Amazon S3 implement de-duplication to identify and remove duplicate files, optimizing user storage space [90]. Similarly, entities such as credit bureaus and e-commerce platforms employ de-duplication methods to consolidate and accurately resolve records based on shared attributes, including names, birth dates, addresses, and transaction histories, even when slight variations exist. Additionally, in the domain of generative AI, training LLMs often involves extensive data de-duplication to improve computational efficiency, enhance model performance, and reduce the risk of privacy leakage by removing redundant or duplicated content from training datasets [91, 92, 93, 94].
In our experiments, we assume that the ML service model is honest-but-curious; if the model does not always provide correct answers to the inference queries, an increased number of queries and higher cost would be incurred.
Currently, LDPKiT only supports the image modality, where superimposition intuitively translates to pixel-level operations. Although textual data does not have a counterpart for superimposition, several conceptual approaches can be explored. These include embedding-level averaging, in which two textual embeddings are blended; token-level mixing, which involves the interleaving of tokens from two samples; and semantic-level merging, such as combining two sentences into a coherent paraphrased summary. Each interpretation may require a distinct adaptation of the LDP scheme. We leave the exploration of these textual superimposition methods as a direction for future research.
As discussed in Section IV-D, LDPKiT is compatible with advanced training strategies, such as active learning, which expedite model training and reduce the required training set size. Active learning strategies can also be applied to prune queries in $\mathcal{D}_{\rm infer}$} that are less significant. We plan to study the effect of such strategies on LDPKiT in the future. Furthermore, since $\mathcal{M}_{\rm L}$} is trained on noisy data, LDPKiT is inherently immune to membership inference attacks if $\mathcal{M}_{\rm L}$} is ever leaked. We leave the analyses of membership inference and attribute inference attack success rates as future work.
Further, our privacy-preserving queries in this paper refer to LDPKiT-generated noised queries with LDPKiT-Rand or LDPKiT-Sup utilizing $\epsilon$-LDP privacy mechanism. We can generate such queries in other ways. For instance, LDPKiT is compatible with other noise mechanisms such as ($\epsilon$, $\delta$)-LDP with the Gaussian mechanism [32]. In addition, we can also compose our privacy-preserving $\mathcal{D}_{\rm priv}$} by generating synthetic queries using a generative model such as a diffusion model or an LLM. We leave a detailed analysis of these alternate techniques for future work.
We also extended the evaluation of LDPKiT to the ImageNet1K dataset [95], where all images were resized to 256×256 pixels and subsequently center-cropped to 224×224 pixels, consistent with the standard preprocessing procedure in [96]. For reference, the official torchvision ResNet-18 model reports a top-1 accuracy of 69.758% on ImageNet1K [96]. In our evaluation, a randomly selected subset of 15 classes was used. Under this setting, LDPKiT-Sup with ResNet-18 ($\mathcal{M}_{\rm L}$}) achieved a top-1 accuracy of 63% with 15,000 $\mathcal{D}_{\rm priv}$} samples at $\epsilon=5$, whereas the baseline SIDP attained only 19% accuracy under identical conditions. Further enhancing the performance of LDPKiT on larger datasets is left for future work.
VII Conclusion
LDPKiT provides a novel privacy-preserving inference framework designed to protect sensitive user data when interacting with powerful yet potentially malicious ML model providers. By incorporating LDP noise into inference queries, LDPKiT protects privacy even if service platforms or models are compromised. The key contribution of LDPKiT is the integration of LDP with a two-layer data augmentation mechanism (LDPKiT-Rand and LDPKiT-Sup), leveraging LDP’s post-processing property to enhance utility without compromising privacy. The first layer provides the baseline LDP privacy protection, and the second layer recovers utility with a larger augmented dataset generated for effective local model extraction. Specifically, LDPKiT-Sup employs a novel superimposition-based augmentation technique to produce more representative data points that better approximate the original private data distribution compared to the random-noise strategy in LDPKiT-Rand. The extensive evaluation results show that LDPKiT-Sup successfully recovers prediction accuracy on private data while preserving privacy. It has greater benefits at higher noise levels, corresponding to stronger privacy guarantees. Furthermore, LDPKiT explicitly acknowledges the economic and ethical considerations: it preserves service providers’ revenue models with user-issued queries and ensures that the extracted models are non-stealthy and non-competitive, complying with providers’ terms of use. In doing so, LDPKiT maintains an ethical and secure design that upholds the principles of responsible data usage and privacy by design.
VIII LLM Usage Considerations
The authors used generative AI-based tools, including OpenAI’s ChatGPT [97] and Grammarly [98], to correct typos, grammatical errors, and awkward phrasings in the preparation of this manuscript. All AI-generated outputs were inspected by the authors to ensure accuracy and originality.
Appendix A Appendix
(a) Label 2
(b) Label 1
Figure 8: Samples of noised SVHN data with $\epsilon$ set to 1.5 (left) and 1.25 (right).
(a) Label 0 (T-shirt/top)
(b) Label 6 (shirt)
Figure 9: Samples of noised Fashion-MNIST data with $\epsilon$ set to 2.0 (left) and 1.5 (right).
(a) Label 0 (adipose)
(b) Label 0 (adipose)
Figure 10: Samples of noised PathMNIST data with $\epsilon$ set to 10.0 (left) and 7.0 (right).
A-A Proof of $\epsilon$-LDP with the Laplacian mechanism and Post-Processing Property
We discuss the proof aforementioned in Section III here: We prove that the base noise injection scheme (i.e., the first layer of noise) generating $\mathcal{D}_{\rm protected}$} satisfies the definition of $\epsilon$}-LDP with the Laplacian mechanism in Theorem A.1.
Theorem A.1.
Our base noise injection mechanism satisfies $\epsilon$}-LDP where $\epsilon=\frac{\Delta_{f}}{\lambda}$
Proof.
Let $v$ be the original data and $f(v)$ be the query and computation function performed on the data with function sensitivity, $\Delta_{f}=\max_{v_{1},v_{2}}\|f(v_{2})-f(v_{1})\|_{1}$. We define the randomization algorithm $\mathcal{A}$ with Laplacian mechanism such that for any input value $v$, $\mathcal{A}=f(v)+\mathcal{Z}$, where $\mathcal{Z}$ is sampled from the Laplacian distribution $\mathcal{L}(z,\lambda)$ with scaling factor $\lambda$ set to $\frac{\Delta_{f}}{\epsilon}$.
The probability that $\mathcal{A}$ has an output $s\in\mathcal{S}$ given an input $v$ can be expressed as:
$\Pr[\mathcal{A}(v)=s]$ $=\Pr[f(v)+\mathcal{Z}=s]$
$=\Pr[\mathcal{Z}=s-f(v)]$
$=\frac{1}{2\lambda}\exp\left(-\frac{|s-f(v)|}{\lambda}\right)$
for all $s\in\mathcal{S}$.
To satisfy $\epsilon$}-LDP, we need Equation 1 to hold for any output $s\in\mathcal{S}$ and any input pairs $v_{1}$ and $v_{2}$.
By substituting the PDF of the Laplacian distribution’s new expression into Equation 1, we get
$\frac{1}{2\lambda}\exp\left(-\frac{|s-f(v_{1})|}{\lambda}\right)\leq e^{\epsilon}\frac{1}{2\lambda}\exp\left(-\frac{|s-f(v_{2})|}{\lambda}\right)$
which simplifies to
$\exp\left(\frac{|s-f(v_{2})|-|s-f(v_{1})|}{\lambda}\right)\leq e^{\epsilon}$
With $\lambda=\frac{\Delta_{f}}{\epsilon}$, the equation becomes:
$\exp\left(\epsilon\cdot\frac{\left|s-f(v_{1})\right|-\left|s-f(v_{2})\right|}{\Delta_{f}}\right)\leq e^{\epsilon}$ (4)
for all pairs of $v_{1}$ and $v_{2}$.
Since $\left|s-f(v_{1})\right|-\left|s-f(v_{2})\right|\leq\Delta_{f}$ by the definition of sensitivity function, Equation 4 always holds.
Then, we prove that our data augmentation algorithms, LDPKiT-Rand and LDPKiT-Sup, satisfy the post-processing property of LDP and hence provide the same level of $\epsilon$-LDP privacy guarantee on $\mathcal{D}_{\rm infer}$} as $\mathcal{D}_{\rm protected}$}. For convenience, we assume the dataset used for querying, $\mathcal{D}_{\rm infer}$}, is the maximum possible augmented inference dataset generated from $\mathcal{D}_{\rm priv}$} without post-selection per definitions in Section III-B.
Theorem A.2.
Our data augmentation mechanism LDPKiT-Rand satisfies $\epsilon$}-LDP.
Proof.
For each data point $d_{i}\in\mathcal{D}_{\text{priv}}$, where $i=1,2,\dots,|{\mathcal{D}_{\rm priv}}|$ (note that $|{\mathcal{D}_{\rm protected}}|=|{\mathcal{D}_{\rm priv}}|$), apply the Laplacian mechanism (See Definition II.2) to obtain the first noise-added data point:
$\tilde{d}_{i}^{(1)}=d_{i}+\eta_{i}^{(1)},$
where $\eta_{i}^{(1)}$ is a random variable drawn from the Laplacian distribution with scaling factor $\lambda=\frac{\Delta_{f}}{\epsilon}$.
By Definition II.1, this initial mechanism satisfies $\epsilon$-LDP.
As for the post-processing step in LDPKiT-Rand, for each $\tilde{d}_{i}^{(1)}$, we generate $|{\mathcal{D}_{\rm protected}}|-1$ distinct random versions by adding an additional layer of Laplacian noise with the same scale:
$\tilde{d}_{i,j}^{(2)}=\tilde{d}_{i}^{(1)}+\eta_{j}^{(2)},$
where $i=1,2,\dots,|{\mathcal{D}_{\rm protected}}|,j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|-1$ and $\eta_{j}^{(2)}$ are independent random variables drawn from the Laplacian distribution with scaling factor $\lambda=\frac{\Delta_{f}}{\epsilon}$.
This step is considered post-processing because it operates on $\tilde{d}_{i}^{(1)}$, which is already an output of an $\epsilon$-LDP mechanism $\mathcal{A}$, and does not access the original data $d_{i}$.
Hence, the protected inference dataset $\mathcal{D}_{\rm infer}$} at most consists of all $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ data points $\tilde{d}_{i,j}^{(2)}$:
${\mathcal{D}_{\rm infer}}=\left\{\tilde{d}_{i,j}^{(2)}\;\middle|\;\begin{array}[]{l}i=1,2,\dots,|{\mathcal{D}_{\rm protected}}|;\\ j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|-1\end{array}\right\}.$
By Definition II.3, since the initial mechanism $\mathcal{A}$ satisfies $\epsilon$-LDP, any function $g$ that processes its output without accessing the original data preserves the $\epsilon$-LDP guarantee. The addition of Laplacian noise to $\tilde{d}_{i}^{(1)}$ is a randomized function $g$ that depends only on $\tilde{d}_{i}^{(1)}$ and independent random noise $\eta_{j}^{(2)}$. Hence, the composite mechanism $g(\mathcal{A}(d_{i}))$ satisfies:
$\Pr\left[g(\mathcal{A}(d_{1}))\in\mathcal{T}\right]\leq e^{\epsilon}\Pr\left[g(\mathcal{A}(d_{2}))\in\mathcal{T}\right],$ (5)
for all values $d_{1},d_{2}$ and all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$.
Therefore, each data point $\tilde{d}_{i,j}^{(2)}$ in $\mathcal{D}_{\rm infer}$} satisfies $\epsilon$-LDP due to the post-processing property. The privacy guarantee from the initial noise addition is preserved and the additional noise does not compromise the privacy level.
Theorem A.3.
Our data augmentation mechanism LDPKiT-Sup satisfies $\epsilon$}-LDP.
Proof.
The base noise addition mechanism $\mathcal{A}$ applies the Laplacian mechanism (See Definition II.2) to obtain the noise-added data point:
$\tilde{d}_{i}=d_{i}+\eta_{i},$
where $\eta_{i}$ is a random variable drawn from the Laplacian distribution with probability density function (PDF):
$L(z,\lambda)=\frac{1}{2\lambda}\exp\left(-\frac{|z|}{\lambda}\right),$
and the scaling factor $\lambda$ is set to $\lambda=\frac{\Delta_{f}}{\epsilon}$.
By Definition II.1, the base noise addition mechanism that outputs $\tilde{d}_{i}$ satisfies $\epsilon$-LDP.
We define a post-processing function $g$ that takes the set of noise-added data points $\{\tilde{d}_{1},\tilde{d}_{2},\dots,\tilde{d}_{|}{\mathcal{D}_{\rm protected}}_{|}\}$ and outputs the average of all possible ordered pairs:
${\mathcal{D}_{\rm infer}}=g\left(\{\tilde{d}_{i}\}\right)=\left\{\begin{aligned} &avg(\tilde{d}_{i},\tilde{d}_{j})\;\mid\;i\neq j,\\ &1\leq i,j\leq|{\mathcal{D}_{\rm protected}}|\end{aligned}\right\}.$
The size of $\mathcal{D}_{\rm infer}$} is $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$.
By Definition II.3, if the base noise addition mechanism $\mathcal{A}$ satisfies $\epsilon$-LDP, then any function $g$ applied to its output preserves the $\epsilon$-LDP guarantee.
Since $g$ operates on $\tilde{d}_{i}$ and does not access the original data $d_{i}$, the composite mechanism $g(\mathcal{A}(d_{i}))$ satisfies:
$\Pr\left[g(\mathcal{A}(d_{1}))\in\mathcal{T}\right]\leq e^{\epsilon}\Pr\left[g(\mathcal{A}(d_{2}))\in\mathcal{T}\right],$ (6)
for all values $d_{1},d_{2}$ and all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$.
Similarly, $g$ does not access the original data $d_{j}$ since it has base noise that satisfies $\epsilon$-LDP added (i.e., $\tilde{d}_{j}$) before composition with $\tilde{d}_{i}$. Hence, the privacy guarantee from the initial noise addition in $d_{j}$ is also not compromised.
Therefore, each data point in $\mathcal{D}_{\rm infer}$} satisfies $\epsilon$-LDP. The privacy guarantee from the base layer of noise addition is preserved in $\mathcal{D}_{\rm infer}$} due to the post-processing property.
TABLE V: Final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of SVHN.
Model $\epsilon$
Accuracy on $\mathcal{D}_{\rm priv}$} (%)
Accuracy on $\mathcal{D}_{\rm val}$} (%)
SIDP LDPKiT-Rand LDPKiT-Sup SIDP LDPKiT-Rand LDPKiT-Sup
ResNet-18 30.0 93.8944 ($\pm$0.7050) 79.4593 ($\pm$4.0448) 93.9556 ($\pm$0.8406) 95.8538 ($\pm$0.0596) 74.6765 ($\pm$4.3553) 95.4970 ($\pm$0.1720)
15.0 89.4130 ($\pm$0.6467) 90.1259 ($\pm$0.8296) 93.9407 ($\pm$0.6916) 92.7431 ($\pm$0.0944) 89.3243 ($\pm$1.1354) 95.3587 ($\pm$0.1559)
2.0 21.0657 ($\pm$0.9500) 84.8593 ($\pm$1.7809) 84.6296 ($\pm$1.1161) 29.9870 ($\pm$0.3446) 88.0190 ($\pm$1.4606) 89.1996 ($\pm$0.5978)
1.5 14.1648 ($\pm$0.8290) 69.6148 ($\pm$1.3020) 84.0815 ($\pm$1.2100) 22.7788 ($\pm$0.3117) 75.9971 ($\pm$1.3721) 88.7301 ($\pm$1.0922)
1.25 11.8722 ($\pm$0.6702) 54.8593 ($\pm$3.9160) 82.6741 ($\pm$0.8038) 14.8386 ($\pm$0.2211) 63.0220 ($\pm$4.2599) 87.1334 ($\pm$0.7090)
MobileNetV2 30.0 93.8944 ($\pm$0.7050) 59.8370 ($\pm$2.0202) 93.5407 ($\pm$0.9009) 95.8538 ($\pm$0.0596) 57.5527 ($\pm$2.3437) 94.9067 ($\pm$0.1872)
15.0 89.4130 ($\pm$0.6467) 79.7630 ($\pm$2.4033) 92.9185 ($\pm$0.7444) 92.7431 ($\pm$0.0944) 78.0475 ($\pm$1.7365) 94.5682 ($\pm$0.1779)
2.0 21.0657 ($\pm$0.9500) 78.9037 ($\pm$1.3885) 80.7481 ($\pm$1.4278) 29.9870 ($\pm$0.3446) 83.5886 ($\pm$1.4901) 86.4235 ($\pm$0.9328)
1.5 14.1648 ($\pm$0.8290) 62.5185 ($\pm$3.4171) 79.1481 ($\pm$1.0790) 22.7788 ($\pm$0.3117) 70.0202 ($\pm$2.6289) 85.2818 ($\pm$0.8650)
1.25 11.8722 ($\pm$0.6702) 52.6741 ($\pm$3.6764) 78.8222 ($\pm$1.2728) 14.8386 ($\pm$0.2211) 60.4773 ($\pm$3.6285) 84.2723 ($\pm$0.9145)
SIDP is the inference accuracy of ResNet-152 ($\mathcal{M}_{\rm R}$}) on $\mathcal{D}_{\rm protected}$} without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
TABLE VI: Final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of Fashion-MNIST.
Model $\epsilon$
Accuracy on $\mathcal{D}_{\rm priv}$} (%)
Accuracy on $\mathcal{D}_{\rm val}$} (%)
SIDP LDPKiT-Rand LDPKiT-Sup SIDP LDPKiT-Rand LDPKiT-Sup
ResNet-18 30.0 91.9359 ($\pm$0.3525) 91.6444 ($\pm$0.6625) 93.0222 ($\pm$0.4773) 91.9044 ($\pm$0.1773) 84.0933 ($\pm$1.5904) 92.1978 ($\pm$0.2630)
15.0 89.0204 ($\pm$0.6521) 91.7185 ($\pm$0.5742) 92.8370 ($\pm$0.3182) 89.0178 ($\pm$0.1862) 89.1944 ($\pm$1.1282) 92.2111 ($\pm$0.2654)
2.0 28.6944 ($\pm$1.0409) 68.5185 ($\pm$2.4922) 81.7778 ($\pm$2.7049) 28.5200 ($\pm$0.3539) 68.8311 ($\pm$2.7650) 81.6011 ($\pm$2.4087)
1.5 23.3074 ($\pm$0.8905) 58.0741 ($\pm$4.7133) 73.5037 ($\pm$2.3786) 22.8678 ($\pm$0.1377) 57.9644 ($\pm$4.8297) 73.6200 ($\pm$2.3208)
MobileNetV2 30.0 91.9359 ($\pm$0.3525) 91.6741 ($\pm$0.5390) 92.9750 ($\pm$0.3837) 91.9044 ($\pm$0.1773) 85.5178 ($\pm$1.0833) 92.2800 ($\pm$0.0748)
15.0 89.0204 ($\pm$0.6521) 89.9556 ($\pm$1.8921) 92.6074 ($\pm$0.4440) 89.0178 ($\pm$0.1862) 86.3111 ($\pm$2.7241) 92.1633 ($\pm$0.1699)
2.0 28.6944 ($\pm$1.0409) 59.7778 ($\pm$1.9402) 74.3407 ($\pm$2.7658) 28.5200 ($\pm$0.3539) 59.9878 ($\pm$2.0344) 74.0622 ($\pm$3.1608)
1.5 23.3074 ($\pm$0.8905) 52.0296 ($\pm$4.9529) 63.3852 ($\pm$5.5650) 22.8678 ($\pm$0.1377) 52.3267 ($\pm$4.9828) 63.5922 ($\pm$5.2323)
SIDP is the inference accuracy of ResNet-152 ($\mathcal{M}_{\rm R}$}) on $\mathcal{D}_{\rm protected}$} without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
TABLE VII: Final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of PathMNIST.
Model $\epsilon$
Accuracy on $\mathcal{D}_{\rm priv}$} (%)
Accuracy on $\mathcal{D}_{\rm val}$} (%)
SIDP LDPKiT-Rand LDPKiT-Sup SIDP LDPKiT-Rand LDPKiT-Sup
ResNet-18 30.0 77.6074 ($\pm$0.7298) 84.1333 ($\pm$1.0050) 83.2000 ($\pm$2.5392) 73.9198 ($\pm$0.2818) 78.0718 ($\pm$0.8531) 76.7471 ($\pm$4.6596)
15.0 47.4315 ($\pm$3.3685) 64.0519 ($\pm$4.6521) 81.5704 ($\pm$2.1463) 46.4268 ($\pm$0.1964) 69.1489 ($\pm$3.5085) 76.7193 ($\pm$1.8324)
10.0 28.7444 ($\pm$2.9476) 43.9037 ($\pm$3.9903) 82.11 ($\pm$1.8628) 25.8774 ($\pm$0.2397) 43.4478 ($\pm$6.2963) 80.2043 ($\pm$1.5305)
7.0 22.1759 ($\pm$2.2869) 31.7481 ($\pm$5.6920) 73.29 ($\pm$2.0545) 19.8963 ($\pm$0.2265) 36.2132 ($\pm$6.2276) 76.6543 ($\pm$1.8549)
30.0 77.6074 ($\pm$0.7298) 84.1556 ($\pm$0.6137) 84.2889 ($\pm$0.7318) 73.9198 ($\pm$0.2818) 77.3290 ($\pm$1.2886) 77.9944 ($\pm$1.6459)
MobileNet 15.0 47.4315 ($\pm$3.3685) 51.2519 ($\pm$13.4225) 80.4444 ($\pm$2.2522) 46.4268 ($\pm$0.1964) 51.7394 ($\pm$12.1483) 75.0418 ($\pm$3.5402)
-V2 10.0 28.7444 ($\pm$2.9476) 25.9556 ($\pm$11.1678) 82.8519 ($\pm$1.0540) 25.8774 ($\pm$0.2397) 30.5060 ($\pm$12.7641) 79.9876 ($\pm$1.8408)
7.0 22.1759 ($\pm$2.2869) 21.8519 ($\pm$11.1653) 79.4370 ($\pm$1.4519) 19.8963 ($\pm$0.2265) 31.3138 ($\pm$13.0969) 82.3398 ($\pm$1.3950)
SIDP is the inference accuracy of ResNet-152 ($\mathcal{M}_{\rm R}$}) on $\mathcal{D}_{\rm protected}$} without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
A-B Examples of $\mathcal{D}_{\rm infer}$} data with LDP noise
Figures 8, 9, and 10 present comparisons between the original and noised data samples under different levels of $\epsilon$-LDP noise. They also illustrate the impact of our two privacy-preserving data augmentation mechanisms, LDPKiT-Rand and LDPKiT-Sup.
(a) Fashion-MNIST ($\epsilon=1.5$)
(b) SVHN ($\epsilon=2.0$)
(c) SVHN ($\epsilon=1.25$)
(d) PathMNIST ($\epsilon=7.0$)
Figure 11: Comparison of ResNet18 ($\mathcal{M}_{\rm L}$})’s accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} with $|{\mathcal{D}_{\rm priv}}|=1.5k$ and various $|{\mathcal{D}_{\rm infer}}|$.
Figure 12: Qualitative results of image reconstruction with diffusion models. (Top row: privacy-preserving data generated with LDPKiT-Sup. Bottom row: applying the reconstruction attack on data protected by LDPKiT-Sup).
Figure 13: Sample reconstructed images demonstrating the effectiveness of the denoiser on SVHN and PathMNIST when little noise is added (e.g., $\epsilon$ = 15.0 and 20.0).
A-C Hyperparameter choices and dataset preparation
In this section, we document hyperparameter choices and dataset splits for the experiments in Section IV. We use three image benchmarks for evaluation: SVHN with 10 classes of 32x32x3 cropped street view house number images, Fashion-MNIST with 10 classes of 28x28 greyscale fashionable clothing images, and PathMNIST with 9 classes of 28x28x3 medical images of pathology. We train $\mathcal{M}_{\rm R}$} with a learning rate of 0.1 for 200 epochs on Fashion-MNIST and SVHN and 10 epochs on PathMNIST. $\mathcal{M}_{\rm R}$} is trained on 35k data points for Fashion-MNIST, 48,257 data points for SVHN and 89,996 data points for PathMNIST. $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} are split from the remaining data points unseen by $\mathcal{M}_{\rm R}$}, where $\mathcal{D}_{\rm priv}$} is used to train and evaluate $\mathcal{M}_{\rm L}$}, and $\mathcal{D}_{\rm val}$} is used for pure model generalizability evaluation. Specifically, CIFAR-10 has 15k data points in the candidate pool of $\mathcal{D}_{\rm priv}$}, Fashion-MNIST and SVHN have 25k, and PathMNIST has 10,004. As for $\mathcal{D}_{\rm val}$}, its size is 10k for both CIFAR-10 and Fashion-MNIST, 26,032 for SVHN, and 7,180 for PathMNIST. $\mathcal{D}_{\rm infer}$} in Section IV-B has a size of $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ (around 2.2M). In Section IV-D, $|{\mathcal{D}_{\rm priv}}|$ is set to 125, 250, 500, 1k and 1.5k, whereas $|{\mathcal{D}_{\rm infer}}|$ is set to 15,500, 62,250, 250k and 500k, respectively. In this section, another set of comparison experiments is done with $|{\mathcal{D}_{\rm priv}}|$ fixed at 1.5k and $\mathcal{D}_{\rm infer}$} has a varying size of 250k, 500k, 1M, and 2M (i.e., all possible superimposition pairs, the same as Section IV-B). $\mathcal{M}_{\rm L}$}’s learning rate is 0.1 for Fashion-MNIST and 0.001 for SVHN and PathMNIST. The number of training epochs is set to 15 across all the datasets in all scenarios. For the latent space analysis, we train the VAE for 200 epochs with a learning rate set to 0.001 for Fashion-MNIST. The number of epochs is set to 300, and the learning rate is 0.0001 for SVHN and PathMNIST.
TABLE VIII: Sensitivity Analysis on $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ with ResNet-18 ($\mathcal{M}_{\rm L}$}) (Continued from Table III).
Dataset $\epsilon$ LDPKiT Accuracy on $\mathcal{D}_{\rm priv}$} (%)
$|{\mathcal{D}_{\rm priv}}|$ 125 250 500 1k 1.5k
$|{\mathcal{D}_{\rm infer}}|$ 15,500 15,500 62,250 15,500 62,250 250k 15,500 62,250 250k 500k 15,500 62,250 250k 500k
Fashion- MNIST 1.5 Rand 15.73 17.78 23.53 17.49 27.37 45.58 14.11 28.48 49.65 56.41 15.89 30.22 51.16 53.51
Sup 18.40 22.27 29.02 21.49 36.91 45.33 22.17 37.16 60.20 65.29 22.34 36.45 58.67 68.66
SVHN 2.0 Rand 8.40 10.40 10.00 10.00 10.40 56.88 11.40 10.90 56.60 74.13 9.83 12.27 62.48 74.69
Sup 10.13 10.13 12.53 10.40 20.13 44.69 10.53 23.07 65.02 71.26 10.04 20.51 66.88 79.94
SVHN 1.25 Rand 9.60 10.80 6.80 10.30 6.80 19.40 4.85 10.40 18.43 36.18 10.05 9.23 20.18 62.48
Sup 9.87 10.27 12.53 9.60 13.27 35.87 10.20 13.50 57.13 65.23 9.76 15.22 57.29 71.90
Path- MNIST 7.0 Rand 20.27 19.69 17.07 20.91 22.63 36.76 20.71 21.24 38.60 32.67 18.33 19.60 31.42 34.86
Sup 30.84 29.02 34.36 30.04 35.38 52.13 28.63 34.82 55.81 65.26 32.02 41.13 59.77 63.99
A-D Additional experimental results
This section is complementary to the evaluation results presented in Section IV.
A-D1 Final accuracy on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$}
Tables V, VI and VII show the tabulated final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of SVHN, Fashion-MNIST and SVHN that SIDP, LDPKiT-Rand and LDPKiT-Sup can achieve, which are also presented as bar graphs in Figure 4 in Section IV-B We record the accuracies at the last epoch of training. The results draw the same conclusion as Section IV that LDPKiT helps offset the accuracy trade-offs brought by LDP noise. The tables also show that LDPKiT can almost always achieve higher inference accuracy than SIDP, but the advantage is less obvious when the least noise is added (i.e., $\epsilon=30$). However, there is barely privacy protection in these cases, as SIDP accuracies are also very high, indicating a significant information leakage to $\mathcal{M}_{\rm R}$}. As more LDP noise is added to $\mathcal{D}_{\rm priv}$} to preserve privacy, the gap in utility that LDPKiT provides over SIDP also increases. Therefore, LDPKiT offers greater benefits in regimes with stronger privacy protection and correspondingly more noise.
A-D2 Additional Dataset Size Sensitivity Analysis
The results in this section are complementary to the discussion in Section IV-D. Figure 11 demonstrates the impact of $|{\mathcal{D}_{\rm infer}}|$ with other $\epsilon$ values on the three image benchmarks. The takeaway aligns with the conclusion drawn from Figure 7 that although larger $\mathcal{D}_{\rm infer}$} helps model training and leads to higher accuracies, LDPKiT can already achieve satisfactory accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} when $|{\mathcal{D}_{\rm infer}}|$ is 500k. Table VIII shows the impact of $|{\mathcal{D}_{\rm infer}}|$ and $|{\mathcal{D}_{\rm priv}}|$ with other $\epsilon$ values. All the additional results align with the conclusions drawn from Table III. The main takeaway aligns with the discussion in Section IV-D. LDPKiT’s accuracies increase as $|{\mathcal{D}_{\rm infer}}|$ increases; however, the improvement becomes minor with larger $\mathcal{D}_{\rm infer}$} and the accuracies are sufficiently high when $|{\mathcal{D}_{\rm infer}}|$ is 500k. Furthermore, LDPKiT is less sensitive to $|{\mathcal{D}_{\rm priv}}|$, so LDPKiT can achieve decent accuracies when the user generates query data with a relatively smaller $\mathcal{D}_{\rm priv}$}.
A-D3 Qualitative Demonstration of Image Reconstruction
Figure 12 is complementary to the quantitative measures of image reconstruction success rate, which shows the reconstructed images from noisy data processed with LDPKiT-Sup when $\epsilon$ is low. The qualitative results also show that the reconstruction attack is not effective on data samples protected by LDPKiT-Sup. To demonstrate the authenticity of the denoiser, we also show the reconstruction results when $\epsilon$ is high (i.e., less noise added) in Figure 13.
(a) (4,4)
(b) (8,4)
(c) (8,8)
Figure 14: Latent space plots of PathMNIST class triplets (C3-lymphocytes, C4-mucus, C8-colorectal adenocarcinoma epithelium) and privacy-protected noisy data clusters generated with LDPKiT-Rand and LDPKiT-Sup ($\epsilon=7.0$).
(a) (1,1)
(b) (1,8)
(c) (2,8)
Figure 15: Latent space plots of SVHN class triplets (C1, C2, C8) and privacy-protected noisy data clusters generated with LDPKiT-Rand and LDPKiT-Sup ($\epsilon=2.0$).
TABLE IX: Euclidean distances between centroids of clusters shown in Figure 14 on PathMNIST.
Figure Strategy Class(es) $d(C_{\text{N}},C_{3})$ $d(C_{\text{N}},C_{4})$ $d(C_{\text{N}},C_{8})$
14(a) Sup (4,4) 1.9453 1.5981 2.5698
Rand 4 1.8724 2.5812 3.3226
14(b) Sup (8,4) 1.4984 2.1860 2.0151
Rand 8 1.6577 3.3427 3.0580
14(c) Sup (8,8) 1.6301 2.4562 1.5323
Rand 8 1.6407 3.3359 3.0413
$d(C_{\text{N}},C_{X})$ is the Euclidean distance between the centroids of the noisy data cluster and the Class X cluster.
TABLE X: Euclidean distances between centroids of clusters shown in Figure 15 on SVHN.
Figure Strategy Class(es) $d(C_{\text{N}},C_{1})$ $d(C_{\text{N}},C_{2})$ $d(C_{\text{N}},C_{8})$
15(a) Sup (1,1) 2.6877 3.1031 3.2987
Rand 1 4.4212 3.0084 3.1381
15(b) Sup (1,8) 2.8854 3.0853 2.9836
Rand 1 4.4242 3.0094 3.1386
15(c) Sup (2,8) 3.8513 2.5689 3.1065
Rand 2 5.2321 3.0173 3.5754
$d(C_{\text{N}},C_{X})$ is the Euclidean distance between the centroids of the noisy data cluster and the Class X cluster.
TABLE XI: Normalized Zest distance results with $l_{1}$ distance metric on $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$}.
$\mathcal{M}_{\rm R}$} $\mathcal{M}_{\rm L}$} Mechanism Fashion-MNIST SVHN PathMNIST
$\epsilon$=2.0
$\epsilon$=1.5
$\epsilon$=1.5
$\epsilon$=1.25
$\epsilon$=10.0
$\epsilon$=7.0
ResNet-152 ResNet-18 LDPKiT-Sup 1.5904 1.7448 1.1977 1.2286 10.2420 10.2548
LDPKiT-Rand 1.9605 2.0126 1.6078 1.6785 10.3160 9.5630
MobileNetV2 LDPKiT-Sup 4.41e+07 8.11e+07 1.2175 1.2595 INF INF
LDPKiT-Rand 431.5359 1.5733 1.3581 1.2679 INF INF
An adversarial model extraction attack is detected if the normalized Zest distance is smaller than 1.0.
TABLE XII: Normalized Zest distance results with $l_{2}$ distance metric on $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$}.
$\mathcal{M}_{\rm R}$} $\mathcal{M}_{\rm L}$} Mechanism Fashion-MNIST SVHN PathMNIST
$\epsilon$=2.0
$\epsilon$=1.5
$\epsilon$=1.5
$\epsilon$=1.25
$\epsilon$=10.0
$\epsilon$=7.0
ResNet-152 ResNet-18 LDPKiT-Sup 1.5760 1.7553 1.1753 1.1899 10.2420 10.2548
LDPKiT-Rand 1.8586 2.0222 1.5914 1.8143 10.3160 9.5630
MobileNetV2 LDPKiT-Sup 4.16E+08 9.12E+08 1.1692 1.1756 INF INF
LDPKiT-Rand 3.60E+03 1.5568 1.4110 1.3631 INF INF
An adversarial model extraction attack is detected if the normalized Zest distance is smaller than 1.0.
TABLE XIII: Normalized Zest distance results with $l_{\infty}$ distance metric on $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$}.
$\mathcal{M}_{\rm R}$} $\mathcal{M}_{\rm L}$} Mechanism Fashion-MNIST SVHN PathMNIST
$\epsilon$=2.0
$\epsilon$=1.5
$\epsilon$=1.5
$\epsilon$=1.25
$\epsilon$=10.0
$\epsilon$=7.0
ResNet-152 ResNet-18 LDPKiT-Sup 1.9396 2.2138 1.1204 1.1532 9.1840 9.3871
LDPKiT-Rand 2.1970 2.1826 1.5332 2.2498 9.2258 9.5491
MobileNetV2 LDPKiT-Sup 1.67E+09 5.86E+09 1.1469 1.1004 INF INF
LDPKiT-Rand 1.04E+04 1.6815 1.5919 1.7202 INF INF
An adversarial model extraction attack is detected if the normalized Zest distance is smaller than 1.0.
TABLE XIV: Estimated costs for image queries in USD. Google Vision uses per-request feature pricing; OpenAI and Azure assume low-detail vision images at 85 tokens/input image and \$5 per 1M input tokens.
# Image Queries Google Vision (\$0.0015/img) OpenAI Vision (85 tokens, \$0.000425/img) Azure OpenAI Vision (85 tokens, \$0.000425/img)
15,500 \$23.25 \$6.59 \$6.59
62,250 \$93.38 \$26.46 \$26.46
250,000 \$375.00 \$106.25 \$106.25
500,000 \$750.00 \$212.50 \$212.50
1,000,000 \$1,500.00 \$425.00 \$425.00
2,200,000 \$3,300.00 \$935.00 \$935.00
A-D4 Additional results of latent space analysis
Apart from the latent space discussion in Section IV-C, we show more analyses of other class triplets and datasets in this section.
Figure 14 presents an example of latent space visualization. Our primary observation is that, regardless of the base image’s class, $\mathcal{D}_{\rm infer}$} generated by LDPKiT-Sup is consistently more accurate compared to LDPKiT-Rand. As demonstrated in Figures 14(a) and 14(c), LDPKiT-Sup consistently forms a cluster whose distribution is closer to the target class distribution than LDPKiT-Rand. Furthermore, Figure 14(b) shows that LDPKiT-Sup’s cluster is positioned between the clusters of Class 4 and Class 8, while LDPKiT-Rand, despite being generated on Class 8 data, remains further from the correct distribution. We validate these observations by calculating the Euclidean distances between cluster centroids. As shown in Table IX, while both LDPKiT-Sup and LDPKiT-Rand clusters are naturally closer to the Class 3 distribution, LDPKiT-Sup can generate a $\mathcal{D}_{\rm infer}$} cluster that aligns more closely with the target class of the base image. For instance, LDPKiT-Sup’s cluster is closest to Class 4 when superimposition is applied to two Class 4 data points. In contrast, LDPKiT-Rand remains relatively static, always closest to the Class 3 distribution regardless of the base image’s class. The results of SVHN in Table X and Figure 15 align with our observations on other datasets. Visually, LDPKiT-Rand’s distribution remains relatively more static than LDPKiT-Sup, regardless of the label of the base image. LDPKiT-Sup, on the other hand, moves closer to the target cluster(s) of the two superimposed images.
A-E Model extraction attack detection with Zest
This section discusses more details of Zest [99]’s usage and results.
A-E1 Steps of model extraction detection
According to the Zest paper [99], we detect model extraction in the following steps:
  1. 1.
    Calculate the Zest distance between the two models to compare, i.e., $D_{z}({\mathcal{M}_{\rm R}},{\mathcal{M}_{\rm L}})$, where $\mathcal{M}_{\rm L}$} is trained on the entire data split of the noisy $\mathcal{D}_{\rm priv}$}, disjoint from $\mathcal{M}_{\rm R}$}’s training dataset.
  2. 2.
    Calculate a reference distance by computing the average distance between five pairs of the victim and extracted models, denoted as $\mathcal{M}_{\rm V}$ and $\mathcal{M}_{\rm E}$, where $\mathcal{M}_{\rm E}$ are generated by training on $\mathcal{M}_{\rm V}$’s labeled training dataset, i.e.,
    $D_{ref}=\frac{1}{5}\sum_{i=1}^{5}D_{z}(\mathcal{M}_{\rm Vi},\mathcal{M}_{\rm Ei}).$
    Here, $\mathcal{M}_{\rm V}$ has the same model architecture as $\mathcal{M}_{\rm R}$}, and $\mathcal{M}_{\rm E}$ has the same model architecture as $\mathcal{M}_{\rm L}$}, but trained on the same dataset as $\mathcal{M}_{\rm V}$, rather than the noisy $\mathcal{D}_{\rm priv}$}.
  3. 3.
    Calculate the normalized Zest distance, i.e., $\overline{D_{z}}=\frac{D_{z}}{D_{ref}}$.
  4. 4.
    Determine the existence of model extraction by comparing $\overline{D_{z}}$ with threshold 1.0 (as set in [99]). $\overline{D_{z}}<1.0$ indicates $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$} are similar models and model extraction exists.
    $\overline{D_{z}}>1.0$ indicates $\mathcal{M}_{\rm R}$} and $\mathcal{M}_{\rm L}$} are dissimilar and model extraction attack does not exist.
A-E2 Zest distances supplementary results
We show the Cosine Distance measurement in the main paper as it is claimed to be the most accurate metric [99]. We provide additional experimental results with $l_{1}$, $l_{2}$ and $l_{\infty}$ distance metrics supported by Zest in Tables XIXII and XIII on $\mathcal{M}_{\rm L}$} generated from Section IV-B. None of the extraction performed in LDPKiT demonstrates a Zest distance <1.0<1.0. Therefore, the results draw the same conclusion as Section VI-A that the type of knowledge distillation in LDPKiT does not construct an adversarial model extraction attack quantitatively.
A-F Cost Analysis of LDPKiT on Major Service Providers
Based on the pricing information provided by Google [100], OpenAI [101] and Azure [102]. We estimate the cost of LDPKiT assuming the user makes $|{\mathcal{D}_{\rm infer}}|$ number of queries shown in Table III and Figure 7, ranging from 15,500 to 2.2M. The cost analysis is shown in Table XIV.